OpenStudy (anonymous):

check my answer? Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). i got y = 4/3x + 8

OpenStudy (anonymous):

and 4 = 4/3*-3 + 8 4 = -4 + 8 4 = 4

OpenStudy (anonymous):

oh, so i am right, lol